# Free Descriptions of Convex Sets

### v2

@@ 3 Structural Results on Free Descriptions
It is straightforward to check that both sequences are freely-described. In particular, all the relevant spaces of invariants stabilize starting at $n=2$, so that the extension of either description of $[-1,1]^{2}$ to a freely-described sequence is unique. However, the two sequences of sets have notably different properties. The first sequence is both intersection and projection compatible while the second is neither. Furthermore, note that $\mathcal{C}_{2}^{(1)}=\mathcal{C}_{2}^{(2)}=[-1,1]^{2}$ but $\mathcal{C}_{n}^{(1)}\neq\mathcal{C}_{n}^{(2)}$ for all $n\neq 2$. Thus, the extension of the unit square $[-1,1]^{2}$ to a freely-described sequence is not unique, but depends on its conic description.
Motivated by this example, we give conditions on free descriptions ensuring that the corresponding sequence of sets satisfies compatibility conditions. In turn, these yield conditions on a description of a fixed set ensuring that it extends to a freely-described and compatible sequence of sets as well as conditions on the existence of limiting descriptions.
3.1 Free Descriptions Certifying Compatibility
Example 3.1 shows that a freely-described sequence of convex sets need not satisfy either compatibility condition in Definition 1.8. In this section, we therefore give conditions on free descriptions ensuring that our compatibility conditions are satisfied.
@@ 3.1 Free Descriptions Certifying Compatibility
  • Free descriptions certifying compatibility often extend to descriptions of infinite-dimensional limits, see Theorem 3.6 below.
@@ 3.2 Extending a Convex Set to a Freely-Described Sequence
Let $\mathscr{V}=\{\mathbb{V}_{n}\}$ be a consistent sequence of $\{\mathsf{G}_{n}\}$-representations. In this section, we fix $n_{0}$ and seek to extend a convex subset $\mathcal{C}_{n_{0}}\subseteq\mathbb{V}_{n_{0}}$ to a freely-described and compatible sequence. As we saw in Example 3.1, different extensions may be obtained from different fixed-dimensional conic descriptions of $\mathcal{C}_{n_{0}}$. We therefore further fix consistent sequences $\mathscr{U}=\{\mathbb{U}_{n}\},\mathscr{W}=\{\mathbb{W}_{n}\}$ and cones $\{\mathcal{K}_{n}\subseteq\mathbb{U}_{n}\}$ satisfying both intersection and projection compatibility, and assume $\mathcal{C}_{n_{0}}$ is described by (ConicSeq). We now ask when the description of $\mathcal{C}_{n_{0}}$ can be extended to a free description that yields a compatible sequence of convex sets.
@@
Now suppose =$\mathscr{F}=\{\mathbb{F}_{n}\}$ is an algebraically free $\mathscr{V}$-module as above with a surjection $\mathscr{F}\to\mathscr{V}$ whose kernel $\mathscr{K}=\{\mathbb{K}_{n}\}$ is generated in degree $k$. Define the composition
@@
In Section 6, we use Theorem 3.5 to computationally parametrize and search over free descriptions certifying compatibility.
@@ 3.3 Limits of Freely-Described Convex Sets
Our last structural result gives conditions under which free descriptions extend to descriptions of continuous limits of convex sets. The certificates of compatibility in Proposition 3.2 play a major role once again in the existence of these limiting descriptions. If $\mathscr{C}=\{\mathcal{C}_{n}\subseteq\mathbb{V}_{n}\}$ is an intersection-compatible sequence of convex subsets of a consistent sequence $\{\mathbb{V}_{n}\}$, define $\mathcal{C}_{\infty}=\bigcup_{n}\mathcal{C}_{n}$, which is a convex subset of $\mathbb{V}_{\infty}$.
@@ 4.1 Simplices and Norms
[…]
@@
In contrast, consider the sequence of solid simplices $\Delta^{n}_{s}=\{x\in\mathbb{R}^{n}:x\geq 0,\mathbbm{1}_{n}^{\top}x\leq 1\}$. This sequence of conic descriptions is the same as above except that here $\mathcal{K}_{n}=\mathbb{R}^{n+1}_{+}$, and in particular is free and certifies intersection compatibility of $\{\Delta^{n}_{s}\}$. However, the sequence $\{\Delta^{n}_{s}\}$ is also projection-compatible, but the above free description of it does not certify this compatibility. That is because $\{A_{n}^{*}\}$ is not a morphism, hence the hypotheses of the second part of Proposition 3.2(a) are not satisfied. Indeed, the above description of the solid simplex does not make its projection compatibility apparent, since $\mathbbm{1}_{n}^{\top}\mathcal{P}_{\mathbb{V}_{n}}x$ can be arbitrary compared to $\mathbbm{1}_{n+1}^{\top}x$ for general $x\in\mathbb{R}^{n+1}$, but must be smaller if $x\geq 0$.
Instead, the following is a free semidefinite description of the solid simplices that certifies both intersection and projection compatibilities:
@@
which is of the form (ConicSeq) with $\mathscr{U}=\mathrm{Sym}^{2}(\mathrm{Sym}^{\leq 1}\mathscr{V})=\{\mathbb{S}^{n+1}\}$, $\mathscr{K}=\{\mathbb{S}^{n+1}_{+}\}$, and $\mathscr{W}=0$. Here both $\{A_{n}\}$ and $\{A_{n}^{*}\}$ are morphisms. It would be interesting to find a free linear programming description of the solid simplices which certifies both compatibilities, or to show that none exists.
@@
Several related free descriptions are derived from the above. The sequence of $\ell_{1}$ unit balls $\mathcal{B}_{\ell_{1}}^{n}=\{x\in\mathbb{R}^{n}:\|x\|_{1}\leq 1\}$ is intersection and projection compatible. It can be written as $\mathcal{B}_{\ell_{1}}^{n}=\{x\in\mathbb{R}^{n}:\exists y\in\Delta^{n}_{s}\textrm{ s.t. }-y\leq x\leq y\}$, which combined with (9) yields a free description certifying both compatibilities. Similarly, the sequence of $\ell_{2}$ unit balls $\mathcal{B}_{\ell_{2}}^{n}=\{x\in\mathbb{R}^{n}:\|x\|_{2}\leq 1\}$ is intersection and projection compatible, and its standard SDP description $\mathcal{B}_{\ell_{2}}^{n}=\left\{x\in\mathbb{R}^{n}:\begin{bmatrix}1&x^{\top}\\ x&I_{n}\end{bmatrix}=A_{n}x+u_{n}\succeq 0\right\}$ is free and certifies these compatibilities similarly to (9).
@@ 4.2 Regular Polygons
Because $\mathbb{V}_{n}=\mathbb{V}_{n+1}$ while $\mathcal{C}_{n}\neq\mathcal{C}_{n+1}$, the sequence $\mathscr{C}$ satisfies neither intersection nor projection compatibility. Nevertheless, it admits the free description
@@
where $\mathscr{W}=\{\mathbb{R}^{2^{n}}\}_{n}$ with embeddings $y\mapsto y\otimes[1,0]^{\top}$, and $\mathscr{U}=\mathscr{V}\oplus\mathscr{W}\oplus\mathbb{R}$. We put the standard inner products on $\mathbb{R}^{2^{n}}$. The group permutes the $2^{n}$ vertices of $\mathcal{C}_{n}$, defining a permutation action on $[2^{n}]$, and it acts on $\mathbb{R}^{2^{n}}$ by applying these permutations to coordinates.
If $\theta_{i}=\pi(2i+1)/2^{n}$ in (10) instead, we mention that the semidefinite description of $\mathcal{C}_{n}$ given in [40] is also free when $\mathscr{U},\mathscr{W},\mathscr{K}$ are chosen appropriately.
4.3 Permutahedra, Schur-Horn orbitopes, and Limits
@@
where the second equality follows by the Birkhoff-von Neumann theorem. The sequence $\{\mathrm{Perm}_{n}\}$, viewed as subsets of the consistent sequence in Example 1.5 with $\mathsf{G}_{n}=\mathsf{S}_{n}$, is neither intersection- nor projection-compatible. Furthermore, their description (11) is not free because the map $M\mapsto M[1,2,\ldots,n]^{\top}$ is not $\mathsf{G}_{n}$-equivariant. The smaller descriptions of these permutahedra in [67, 68] are also not free because they are not equivariant. However, there is a sequence of permutahedra arising naturally from a limiting perspective that is both intersection- and projection-compatible and whose description certifies this compatibility.
Fix $q,m\in\mathbb{N}$ and a vector $\lambda\in\mathbb{R}^{q}$ with distinct entries, and define $\widetilde{\lambda}=(\lambda_{1},\ldots,\lambda_{1},\ldots,\lambda_{q},\ldots,\lambda_{q})\allowbreak\in\mathbb{R}^{m}$ in which $\lambda_{i}$ appears $m_{i}$ times (so $\sum_{i}m_{i}=m$). Let $\mathsf{G}_{n}=\mathsf{S}_{m2^{n}}$ embedded in $\mathsf{G}_{n+1}$ by sending a $m2^{n}\times m2^{n}$ matrix $g\in\mathsf{G}_{n}$ to $g\otimes I_{2}$. Let $\mathscr{V}=\{\mathbb{R}^{m2^{n}}\}_{n}$ with embeddings $x\mapsto x\otimes\mathbbm{1}_{2}$, the normalized inner product $\langle x,y\rangle=(m2^{n})^{-1}x^{\top}y$, and the standard action of $\mathsf{G}_{n}$. We consider convex hulls of all the vectors in $\mathscr{V}$ containing $\lambda_{i}$ in a fraction $m_{i}/m$ of its entries, which are given using the Birkhoff-von Neumann theorem by
@@
=$=\left\{M\lambda:M\in\mathbb{R}^{m2^{n}\times q}_{+},\ M\mathbbm{1}_{q}=\mathbbm{1}_{m2^{n}},\ M^{\top}\mathbbm{1}_{m2^{n}}=2^{n}[m_{1},\ldots,m_{q}]^{\top}\right\}.$
This is an intersection- and projection-compatible sequence of subsets of $\mathscr{V}$. Moreover, the description (12) is free and certifies intersection and projection compatibility. Indeed, let $\mathscr{W}=\mathscr{V}^{\oplus q}=\{\mathbb{R}^{m2^{n}\times q}\}_{n}$ and $\mathscr{U}=\mathscr{W}\oplus\mathscr{V}^{\oplus 2}\oplus\mathbb{R}^{q}$ containing cones {+$\{\mathbb{R}^{m2^{n}\times q}_{+}\oplus 0\oplus 0\}$. Then (12) is of the form (ConicSeq) with
$A_{n}x=(0,-x,0,0),\quad B_{n}M=(M,\ M\lambda,\ M\mathbbm{1}_{q},\ (m2^{n})^{-1}M^{\top}\mathbbm{1}_{m2^{n}}),$
@@
$(\mathcal{P}_{n}f)(x)=m2^{n}\int_{I_{i}^{(n)}}f\quad\textrm{if }x\in I_{i}^{(n)},$ (13)
hence Jensen’s inequality implies $\|\mathcal{P}_{n}f\|_{L^{p}}\leq\|f\|_{L^{p}}$ for all $p\in[1,\infty]$. Thus, we can identify $\mathbb{V}_{\infty}$ with the space of such step functions, which is contained in $L^{p}([0,1))$ for all $p\in[1,\infty]$, and we can consider the completion of $\mathbb{V}_{\infty}$ with respect to any of the $L^{p}$ norms. If $p<\infty$ we get $\overline{\mathbb{V}_{\infty}}=L^{p}([0,1))$ while if $p=\infty$ then $\overline{\mathbb{V}_{\infty}}$ is the space of functions continuous on $[0,1)\setminus\{m/2^{n}:m,n\in\mathbb{N}\}$ and having left and right limits everywhere [69, §7.6]. Then $\overline{\mathrm{Perm}(\lambda)_{\infty}}$ is the closed convex hull of functions in $\overline{\mathbb{V}_{\infty}}$ that take values $\lambda_{i}$ on a subset of $[0,1]$ of measure $m_{i}/m$. Furthermore, Theorem 3.6 applies to the description in (12) and yields the following dense subset of $\overline{\mathrm{Perm}(\lambda)_{\infty}}$.
@@
=$=\overline{\left\{\sum_{i=1}^{q}\lambda_{i}f_{i}:f_{i}\in\overline{\mathbb{V}_{\infty}},\ f_{i}\geq 0\textrm{ a.e. },\sum_{i=1}^{q}f_{i}=1,\ \int_{[0,1)}f_{i}=\frac{m_{i}}{m}\right\}}.$
@@
Schur-Horn orbitopes:
We consider the matrix analogs of the above permutahedra, which are convex hulls of all matrices with a given spectrum. Let $\mathsf{G}_{n}=\mathsf{O}_{m2^{n}}$ embed in $\mathsf{G}_{n+1}$ by sending a $m2^{n}\times m2^{n}$ matrix $g$ to $g\otimes I_{2}$, let $\mathbb{V}_{n}=\mathbb{S}^{m2^{n}}$ embed in $\mathbb{V}_{n+1}$ by $X\mapsto X\otimes I_{2}$ with normalized inner product $\langle X,Y\rangle=(m2^{n})^{-1}\mathrm{Tr}(XY)$, and finally let $\mathsf{G}_{n}$ act by conjugation on $\mathbb{V}_{n}$. Consider the sequence of Schur-Horn orbitopes [70, Eq. (19)]
@@
=$=\left\{\sum_{i=1}^{q}\lambda_{i}Y_{i}:Y_{1},\ldots,Y_{q}\in\mathbb{V}_{n}\textrm{ s.t. }\sum_{i=1}^{q}Y_{i}=I,\ Y_{i}\succeq 0,\ \mathrm{Tr}(Y_{i})=m_{i}2^{n}\textrm{ for }i=1,\ldots,q\right\},$
which is the matrix analog of (12). This is again a free description certifying both intersection- and projection-compatibility. Indeed, let $\mathscr{W}=\mathscr{V}^{\oplus q}$ and $\mathscr{U}=\mathscr{W}\oplus\mathscr{V}^{\oplus 2}\oplus\mathbb{R}^{q}$ containing the cones $\{\mathcal{K}_{n}=(\mathbb{S}^{m2^{n}}_{+})^{\oplus q}\oplus 0\oplus 0\}$. Then (15) is of the form (ConicSeq) with
$A_{n}X=(0,-X,0,0),\quad B_{n}[Y_{i}]_{i=1}^{q}=\left([Y_{i}]_{i=1}^{q},\ \sum\nolimits_{i}\lambda_{i}Y_{i},\ \sum\nolimits_{i}Y_{i},\ (m2^{n})^{-1}[\mathrm{Tr}(Y_{1}),\ldots,\mathrm{Tr}(Y_{q})]\right),$
@@
Again, we can use Theorem 3.6 to describe an infinite-dimensional limit $\overline{\mathrm{SH}(\lambda)_{\infty}}$ of these Schur-Horn orbitopes. We would like to interpret the elements of $\mathrm{SH}(\lambda)_{\infty}$ as operators with spectral measure $\sum_{i=1}^{q}\frac{m_{i}}{m}\delta_{\lambda_{i}}$, then describe the convex hull of all such operators in $\overline{\mathbb{V}_{\infty}}$. To that end, we complete $\mathbb{V}_{\infty}$ with respect to the operator norm and consider spectral measures with respect to the normalized trace $\tau$ on $\mathbb{V}_{\infty}$, which also extends to the limit. Such spectral measures and associated orbitopes exist in more general algebras. In fact, our framework yields descriptions for Schur-Horn orbitopes, as well as a generalization of the Schur-Horn theorem, to self-adjoint elements in so-called approximately finite-dimensional (AF) algebras. These orbitopes generalize both the permutahedra and Schur-Horn orbitopes in this section. We construct these algebras and apply our framework to the resulting orbitopes in Appendix C. In particular, we obtain the following description for the limit in our present setting, proved in Proposition C.2 of that appendix.
@@
Furthermore, the diagonal map $\mathrm{diag}_{n}\colon\mathbb{S}^{m2^{n}}\to\mathbb{R}^{m2^{n}}$ satisfies $\mathrm{diag}_{n}(\mathrm{SH}(\lambda)_{n})=\mathrm{Perm}(\lambda)_{n}$ by the Schur-Horn theorem. Our limiting descriptions yield the following limiting version of this theorem, proved in greater generality in Proposition C.3 of the above appendix.
@@
Schur-Horn orbitopes are special cases of so-called spectral polyhedra studied in [71]. It would be interesting to identify further examples of freely-described sequences of spectral polyhedra arising in applications and to consider their limits.

### v3

@@ 3 Structural Results on Dimension-Free Descriptions
It is straightforward to check that both sequences are freely-described. In particular, all the relevant spaces of invariants stabilize starting at $n=2$, so that the extension of either description of $1,1]^{2}$ to a freely-described sequence is unique. However, the two sequences of sets have notably different properties. The first sequence is both intersection and projection compatible while the second is neither. Furthermore, note that $\mathcal{C}_{2}^{(1)}=\mathcal{C}_{2}^{(2)}=[-1,1]^{2}$ but $\mathcal{C}_{n}^{(1)}\neq\mathcal{C}_{n}^{(2)}$ for all $n\neq 2$. Thus, the extension of the unit square $1,1]^{2}$ to a freely-described sequence is not unique, but depends on its conic description.
Motivated by this example, we give conditions on dimension-free descriptions ensuring that the corresponding sequence of sets satisfies compatibility conditions. In turn, these yield conditions on a description of a fixed set ensuring that it extends to a freely-described and compatible sequence of sets as well as conditions on the existence of limiting descriptions.
3.1 Dimension-Free Descriptions Certifying Compatibility
Example 3.1 shows that a freely-described sequence of convex sets need not satisfy either compatibility condition in Definition 1.8. In this section, we therefore give conditions on dimension-free descriptions ensuring that our compatibility conditions are satisfied.
@@ 3.1 Dimension-Free Descriptions Certifying Compatibility
  • Dimension-free descriptions certifying compatibility often extend to descriptions of infinite-dimensional limits, see Theorem 3.6 below.
@@ 3.2 Extending a Convex Set to a Freely-Described Sequence
Let $\mathscr{V}=\{\mathbb{V}_{n}\}$ be a consistent sequence of $\{\mathsf{G}_{n}\}$-representations. In this section, we fix $n_{0}$ and seek to extend a convex subset $\mathcal{C}_{n_{0}}\subseteq\mathbb{V}_{n_{0}}$ to a freely-described and compatible sequence. As we saw in Example 3.1, different extensions may be obtained from different fixed-dimensional conic descriptions of $\mathcal{C}_{n_{0}}$. We therefore further fix consistent sequences $\mathscr{U}=\{\mathbb{U}_{n}\},\mathscr{W}=\{\mathbb{W}_{n}\}$ and cones $\{\mathcal{K}_{n}\subseteq\mathbb{U}_{n}\}$ satisfying both intersection and projection compatibility, and assume $\mathcal{C}_{n_{0}}$ is described by (ConicSeq). We now ask when the description of $\mathcal{C}_{n_{0}}$ can be extended to a dimension-free description that yields a compatible sequence of convex sets.
@@
Now suppose $\mathscr{F}=\{\mathbb{F}_{n}\}$ is an algebraically free $\mathscr{V}$-module as above with a surjection $\mathscr{F}\to\mathscr{V}$ whose kernel $\mathscr{K}=\{\mathbb{K}_{n}\}$ is generated in degree $k$. Define the composition
@@
In Section 6, we use Theorem 3.5 to computationally parametrize and search over dimension-free descriptions certifying compatibility.
@@ 3.3 Limits of Freely-Described Convex Sets
Our last structural result gives conditions under which dimension-free descriptions extend to descriptions of continuous limits of convex sets. The certificates of compatibility in Proposition 3.2 play a major role once again in the existence of these limiting descriptions. If $\mathscr{C}=\{\mathcal{C}_{n}\subseteq\mathbb{V}_{n}\}$ is an intersection-compatible sequence of convex subsets of a consistent sequence $\{\mathbb{V}_{n}\}$, define $\mathcal{C}_{\infty}=\bigcup_{n}\mathcal{C}_{n}$, which is a convex subset of $\mathbb{V}_{\infty}$.
@@ 4.1 Simplices and Norms
The sequence of simplices $\Delta^{n-1}=\{x\in\mathbb{R}^{n}:x\geq 0,\ \mathbbm{1}_{n}^{\top}x=1\}$ is freely-described, as the associated description is of the form (ConicSeq) with description spaces $\mathscr{U}=\mathscr{V}\oplus\mathbb{R}=\{\mathbb{U}_{n}=\mathbb{R}^{n+1}\}$, $\mathscr{W}=\{\mathbb{W}_{n}=0\}$ and cones $\mathscr{K}=\{\mathcal{K}_{n}=\mathbb{R}^{n}_{+}\oplus\{0\}\}$, and with $A_{n}=[I_{n},-\mathbbm{1}_{n}]^{\top}$, $B_{n}=0$, and $u_{n}=[0,1]^{\top}$. Moreover, the simplices $\{\Delta^{n-1}\}$ are intersection-compatible (but not projection-compatible); the above description satisfies the hypotheses of Proposition 3.2(a) and hence certifies this compatibility.
@@
In contrast, consider the sequence of solid simplices $\Delta^{n}_{s}=\{x\in\mathbb{R}^{n}:x\geq 0,\mathbbm{1}_{n}^{\top}x\leq 1\}$. This sequence of conic descriptions is the same as above except that here $\mathcal{K}_{n}=\mathbb{R}^{n+1}_{+}$, and in particular is dimension-free and certifies intersection compatibility of $\{\Delta^{n}_{s}\}$. However, the sequence $\{\Delta^{n}_{s}\}$ is also projection-compatible, but the above dimension-free description of it does not certify this compatibility. That is because $\{A_{n}^{*}\}$ is not a morphism, hence the hypotheses of the second part of Proposition 3.2(a) are not satisfied. Indeed, the above description of the solid simplex does not make its projection compatibility apparent, since $\mathbbm{1}_{n}^{\top}\mathcal{P}_{\mathbb{V}_{n}}x$ can be arbitrary compared to $\mathbbm{1}_{n+1}^{\top}x$ for general $x\in\mathbb{R}^{n+1}$, but must be smaller if $x\geq 0$.
Instead, the following is a dimension-free semidefinite description of the solid simplices that certifies both intersection and projection compatibilities:
@@
which is of the form (ConicSeq) with $\mathscr{U}=\mathrm{Sym}^{2}(\mathrm{Sym}^{\leq 1}\mathscr{V})=\{\mathbb{S}^{n+1}\}$, $\mathscr{K}=\{\mathbb{S}^{n+1}_{+}\}$, and $\mathscr{W}=0$. Here both $\{A_{n}\}$ and $\{A_{n}^{*}\}$ are morphisms. It would be interesting to find a dimension-free linear programming description of the solid simplices which certifies both compatibilities, or to show that none exists.
@@
Several related dimension-free descriptions are derived from the above. The sequence of $\ell_{1}$ unit balls $\mathcal{B}_{\ell_{1}}^{n}=\{x\in\mathbb{R}^{n}:\|x\|_{1}\leq 1\}$ is intersection and projection compatible. It can be written as $\mathcal{B}_{\ell_{1}}^{n}=\{x\in\mathbb{R}^{n}:\exists y\in\Delta^{n}_{s}\textrm{ s.t. }-y\leq x\leq y\}$, which combined with (9) yields a dimension-free description certifying both compatibilities. Similarly, the sequence of $\ell_{2}$ unit balls $\mathcal{B}_{\ell_{2}}^{n}=\{x\in\mathbb{R}^{n}:\|x\|_{2}\leq 1\}$ is intersection and projection compatible, and its standard SDP description $\mathcal{B}_{\ell_{2}}^{n}=\left\{x\in\mathbb{R}^{n}:\begin{bmatrix}1&x^{\top}\\ x&I_{n}\end{bmatrix}=A_{n}x+u_{n}\succeq 0\right\}$ is dimension-free and certifies these compatibilities similarly to (9).
@@ 4.2 Regular Polygons
Because $\mathbb{V}_{n}=\mathbb{V}_{n+1}$ while $\mathcal{C}_{n}\neq\mathcal{C}_{n+1}$, the sequence $\mathscr{C}$ satisfies neither intersection nor projection compatibility. Nevertheless, it admits the dimension-free description
@@
where $\mathscr{W}=\{\mathbb{R}^{2^{n}}\}_{n}$ with embeddings $y\mapsto y\otimes[1,0]^{\top}$, and $\mathscr{U}=\mathscr{V}\oplus\mathscr{W}\oplus\mathbb{R}$. We put the standard inner products on $\mathbb{R}^{2^{n}}$. The group permutes the $2^{n}$ vertices of $\mathcal{C}_{n}$, defining a permutation action on $2^{n}]$, and it acts on $\mathbb{R}^{2^{n}}$ by applying these permutations to coordinates.
If $\theta_{i}=\pi(2i+1)/2^{n}$ in (10) instead, we mention that the semidefinite description of $\mathcal{C}_{n}$ given in [40] is also dimension-free when $\mathscr{U},\mathscr{W},\mathscr{K}$ are chosen appropriately.
4.3 Permutahedra, Schur–Horn Orbitopes, and Limits
@@
where the second equality follows by the Birkhoff–von Neumann theorem. The sequence $\{\mathrm{Perm}_{n}\}$, viewed as subsets of the consistent sequence in Example 1.5 with $\mathsf{G}_{n}=\mathsf{S}_{n}$, is neither intersection- nor projection-compatible. Furthermore, their description (11) is not dimension-free because the map $M\mapsto M[1,2,\ldots,n]^{\top}$ is not $\mathsf{G}_{n}$-equivariant. The smaller descriptions of these permutahedra in [65, 66] are also not dimension-free because they are not equivariant. However, there is a sequence of permutahedra arising naturally from a limiting perspective that is both intersection- and projection-compatible and whose description certifies this compatibility.
Fix $q,m\in\mathbb{N}$ and a vector $\lambda\in\mathbb{R}^{q}$ with distinct entries, and define $\widetilde{\lambda}=(\lambda_{1},\ldots,\lambda_{1},\ldots,\lambda_{q},\ldots,\lambda_{q})\allowbreak\in\mathbb{R}^{m}$ in which $\lambda_{i}$ appears $m_{i}$ times (so $\sum_{i}m_{i}=m$). Let $\mathsf{G}_{n}=\mathsf{S}_{m2^{n}}$ embedded in $\mathsf{G}_{n+1}$ by sending a $m2^{n}\times m2^{n}$ matrix $g\in\mathsf{G}_{n}$ to $g\otimes I_{2}$. Let $\mathscr{V}=\{\mathbb{R}^{m2^{n}}\}_{n}$ with embeddings $x\mapsto x\otimes\mathbbm{1}_{2}$, the normalized inner product $\langle x,y\rangle=(m2^{n})^{-1}x^{\top}y$, and the standard action of $\mathsf{G}_{n}$. We consider convex hulls of all the vectors in $\mathscr{V}$ containing $\lambda_{i}$ in a fraction $m_{i}/m$ of its entries, which are given using the Birkhoff–von Neumann theorem by
@@
$=\left\{M\lambda:M\in\mathbb{R}^{m2^{n}\times q}_{+},\ M\mathbbm{1}_{q}=\mathbbm{1}_{m2^{n}},\ M^{\top}\mathbbm{1}_{m2^{n}}=2^{n}[m_{1},\ldots,m_{q}]^{\top}\right\}.$
This is an intersection- and projection-compatible sequence of subsets of $\mathscr{V}$. Moreover, the description (12) is dimension-free and certifies intersection and projection compatibility. Indeed, let $\mathscr{W}=\mathscr{V}^{\oplus q}=\{\mathbb{R}^{m2^{n}\times q}\}_{n}$ and $\mathscr{U}=\mathscr{W}\oplus\mathscr{V}^{\oplus 2}\oplus\mathbb{R}^{q}$ containing cones $\{\mathbb{R}^{m2^{n}\times q}_{+}\oplus 0\oplus 0\}$. Then (12) is of the form (ConicSeq) with
$A_{n}x=(0,-x,0,0),\qquad B_{n}M=(M,\ M\lambda,\ M\mathbbm{1}_{q},\ (m2^{n})^{-1}M^{\top}\mathbbm{1}_{m2^{n}}),$
@@
[…]
hence Jensen’s inequality implies $\|\mathcal{P}_{n}f\|_{L^{p}}\leq\|f\|_{L^{p}}$ for all $p\in[1,\infty]$. Thus, we can identify $\mathbb{V}_{\infty}$ with the space of such step functions, which is contained in $L^{p}([0,1))$ for all $p\in[1,\infty]$, and we can consider the completion of $\mathbb{V}_{\infty}$ with respect to any of the $L^{p}$ norms. If $p<\infty$ we get $\overline{\mathbb{V}_{\infty}}=L^{p}([0,1))$ while if $p=\infty$ then $\overline{\mathbb{V}_{\infty}}$ is the space of functions continuous on $0,1)\setminus\{m/2^{n}:m,n\in\mathbb{N}\}$ and having left and right limits everywhere [67, §7.6]. Then $\overline{\mathrm{Perm}(\lambda)_{\infty}}$ is the closed convex hull of functions in $\overline{\mathbb{V}_{\infty}}$ that take values $\lambda_{i}$ on a subset of $[0,1]$ of measure $m_{i}/m$. Furthermore, Theorem 3.6 applies to the description in (12) and yields the following dense subset of $\overline{\mathrm{Perm}(\lambda)_{\infty}}$.
@@
$=\overline{\left\{\sum_{i=1}^{q}\lambda_{i}f_{i}:f_{i}\in\overline{\mathbb{V}_{\infty}},\ f_{i}\geq 0\textrm{ a.e. },\sum_{i=1}^{q}f_{i}=1,\ \int_{[0,1)}f_{i}=\frac{m_{i}}{m}\right\}}.$
@@
Schur–Horn orbitopes:
We consider the matrix analogs of the above permutahedra, which are convex hulls of all matrices with a given spectrum. Let $\mathsf{G}_{n}=\mathsf{O}_{m2^{n}}$ embed in $\mathsf{G}_{n+1}$ by sending a $m2^{n}\times m2^{n}$ matrix $g$ to $g\otimes I_{2}$, let $\mathbb{V}_{n}=\mathbb{S}^{m2^{n}}$ embed in $\mathbb{V}_{n+1}$ by $X\mapsto X\otimes I_{2}$ with normalized inner product $\langle X,Y\rangle=(m2^{n})^{-1}\mathrm{Tr}(X^{\top}Y)$, and finally let $\mathsf{G}_{n}$ act by conjugation on $\mathbb{V}_{n}$. Consider the sequence of Schur–Horn orbitopes [68, Eq. (19)]
@@
[…]
which is the matrix analog of (12). This is again a dimension-free description certifying both intersection- and projection-compatibility. Indeed, let $\mathscr{W}=\mathscr{V}^{\oplus q}$ and $\mathscr{U}=\mathscr{W}\oplus\mathscr{V}^{\oplus 2}\oplus\mathbb{R}^{q}$ containing the cones $\{\mathcal{K}_{n}=(\mathbb{S}^{m2^{n}}_{+})^{\oplus q}\oplus 0\oplus 0\}$. Then (15) is of the form (ConicSeq) with
$A_{n}X=(0,-X,0,0),\qquad B_{n}[Y_{i}]_{i=1}^{q}=\left([Y_{i}]_{i=1}^{q},\ \sum\nolimits_{i}\lambda_{i}Y_{i},\ \sum\nolimits_{i}Y_{i},\ (m2^{n})^{-1}[\mathrm{Tr}(Y_{1}),\ldots,\mathrm{Tr}(Y_{q})]\right),$
@@
Again, we can use Theorem 3.6 to describe an infinite-dimensional limit $\overline{\mathrm{SH}(\lambda)_{\infty}}$ of these Schur–Horn orbitopes. We would like to interpret the elements of $\mathrm{SH}(\lambda)_{\infty}$ as operators with spectral measure $\sum_{i=1}^{q}\frac{m_{i}}{m}\delta_{\lambda_{i}}$, then describe the convex hull of all such operators in $\overline{\mathbb{V}_{\infty}}$. To that end, we complete $\mathbb{V}_{\infty}$ with respect to the operator norm and consider spectral measures with respect to the normalized trace $\tau$ on $\mathbb{V}_{\infty}$, which also extends to the limit. Such spectral measures and associated orbitopes exist in more general algebras. In fact, our framework yields descriptions for Schur–Horn orbitopes, as well as a generalization of the Schur–Horn theorem, to self-adjoint elements in so-called approximately finite-dimensional (AF) algebras. These orbitopes generalize both the permutahedra and Schur–Horn orbitopes in this section. We construct these algebras and apply our framework to the resulting orbitopes in Appendix C. In particular, we obtain the following description for the limit in our present setting, proved in Proposition C.2 of that appendix.
@@
Furthermore, the diagonal map $\mathrm{diag}_{n}\colon\mathbb{S}^{m2^{n}}\to\mathbb{R}^{m2^{n}}$ satisfies $\mathrm{diag}_{n}(\mathrm{SH}(\lambda)_{n})=\mathrm{Perm}(\lambda)_{n}$ by the Schur–Horn theorem. Our limiting descriptions yield the following limiting version of this theorem, proved in greater generality in Proposition C.3 of the above appendix.
@@
Schur–Horn orbitopes are special cases of so-called spectral polyhedra studied in [69]. It would be interesting to identify further examples of freely-described sequences of spectral polyhedra arising in applications and to consider their limits.